Which equation has a solution of (2,0)
A. 3x - y = 6; y + 2 = x
B. 3x + y = 6; y + 2 = x

Answer by Edwin McCravy(20055)

Which equation has a solution of (2,0)
The way to find out is to plug in (2) for x and (0) for y
the simplify to see if both equations come out with the  
same number on both sides of the equation.

Let's find out if (2,0) is a solution to

A. 3x - y = 6; y + 2 = x
We'll plug in (2) for x and (0) for y:

      3(2) - (0) = 6;  (0) + 2 = (2)
         6 -   0 = 6;        2 = 2
               6 = 6;

Yes, (2,0) is a solution because when we plugged in (2) for x and (0)
for y then simplified, we saw that both equations came out with the 
same number on both sides of the equation.
  
Now Let's find out if (2,0) is a solution to

B. 3x + y = 6; y + 2 = x
We'll plug in (2) for x and (0) for y:

      3(2) + (0) = 6;  (0) + 2 = (2)
         6 +   0 = 6;        2 = 2
               6 = 6;

Yes, (2,0) is a solution because when we plugged in (2) for x and (0)
for y, then simplified, we saw that both equations came out with the 
same number on both sides of the equation.

Edwin